1. An airplane weighs 18,000lbs. The local gravitational acceleration g is 32fps^2. What is the mass of the airplane?

2.The same airplane accelerates down the runway with a net force of 6000lb. What is the acceleration of the airplane?

3.If the same airplane as above airplane starts from a brakes locked position on the runway and it takes off at an airspeed of 200fps. What is the time for the aircraft to reach take-off speed?

4.Under no-wind conditions, what take-off roll is required for the aircraft above?

5.Upon reaching a velocity of 100fps, the pilot of the airplane decides to abort the take off and applies brakes and stops the airplane in 1000ft. What is the airplane’s deceleration? in absolute value
Answer #1

Q1.
Weight = 18000lbs
g = 32fps^2
m = ?
W = m*g
=> m = W/g = 18000/32
=> m = 562.5 slugs

Q2.
F = 6000lb
F = m*a
=> 6000 = 562.5*a
=> a = 10.7 fps^2

Please post other questions separately.